The tens digit of a three-digit number is 0. The sum of the other digits is 6. Interchanged the units and hundreds digits decreases the number by 396. Find the original number
Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

The tens digit of a three-digit number is 0. The sum of the other digits is 6. Interchanged the units and hundreds digits decreases the number by 396. Find the original number
------------------
396/99 = 4, the difference between the 2 digits.
---> 1 and 5
==========================
The tens digit is not relevant.

Answer by greenestamps(13200)



Given the two 3-digit numbers ABC and CBA, and assuming A is greater than C, then the difference ABC-CBA is equal to 99 times (A-C).

Since that difference is 396, we know A-C is 396/99 = 4.

So the hundreds digit is 4 more than the units digit; and the problem says the tens digit is 0 and the sum of the other two digits is 6. Then we have

A-C = 4
A+C = 6

which gives us A=5 and C=1.

ANSWER: The original number is 501